Whatever

This article, which looks like L^{A}TEX is actually an SGML application edited with L^{A}TEX-like input. Document network location:

   DVI: http://www.albany.edu/~hammond/gellmu/mui/tm.dvi
 HTML: http://www.albany.edu/~hammond/gellmu/mui/tm.html

[Processed from GELLMU to L^{A}TEX on Fri Feb 11 08:51:55 EST 2000]